### v1

# Simulation-based inference via telescoping ratio estimation for trawl processes

Abstract
The growing availability of large and complex datasets has increased interest in temporal stochastic processes that can capture stylized facts such as marginal skewness, non-Gaussian tails, long memory, and even non-Markovian dynamics. While such models are often easy to simulate from, parameter estimation remains challenging. Simulation-based inference (SBI) offers a promising way forward, but existing methods typically require large training datasets or complex architectures and frequently yield confidence (credible) regions that fail to attain their nominal values, raising doubts on the reliability of estimates for the very features that motivate the use of these models. To address these challenges, we propose a fast and accurate, sample-efficient SBI framework for amortized posterior inference applicable to intractable stochastic processes. The proposed approach relies on two main steps: first, we learn the posterior density by decomposing it sequentially across parameter dimensions. Then, we use Chebyshev polynomial approximations to efficiently generate independent posterior samples, enabling accurate inference even when Markov chain Monte Carlo methods mix poorly. We further develop novel diagnostic tools for SBI in this context, as well as post-hoc calibration techniques; the latter not only lead to performance improvements of the learnt inferential tool, but also to the ability to reuse it directly with new time series of varying lengths, thus amortizing the training cost. We demonstrate the method’s effectiveness on trawl processes, a class of flexible infinitely divisible models that generalize univariate Gaussian processes, applied to energy demand data.
Keywords: Chebyshev polynomials, Kernel convolution, Neural posterior inference, Neural ratio estimation, Non-Gaussian time series, Non-Markovian processes.
1 Introduction
Gaussian processes (GPs) have become essential tools for modeling and quantifying uncertainty in complex systems, due to their rich mathematical theory and analytic tractability. However, the Gaussianity assumption is often unrealistic. Time series with skewed and heavy-tailed distributions can be observed in macroeconomics, e.g., in financial asset returns (Bradley & Taqqu 2003) or in earth and climate sciences, e.g., in natural phenomena such as earthquakes and floods (Caers et al. 1999), among others. Beyond empirical observations, some physical systems are subject to inherently non-Gaussian noise that standard GPs fail to capture (Vio et al. 2001). While hierarchical models offer one path toward incorporating non-Gaussianity (Rue et al. 2009), they sacrifice the analytical tractability of marginal distributions and autocorrelations, and also face computational challenges when handling high-dimensional covariance matrices. Therefore, it is crucial to develop processes that can directly model non-Gaussianity.
One promising class of non-Gaussian temporal processes is trawl processes, which emerges within the broader framework of Ambit stochastics (Barndorff-Nielsen et al. 2018). This framework provides a general theory for modeling spatio-temporal phenomena and also encompasses certain stochastic partial differential equations (SPDEs) as special cases; see Barndorff-Nielsen et al. (2011) for a detailed discussion on this connection. Trawl processes themselves extend continuous-time GPs to the infinitely divisible setting. These processes, which may be real-valued or integer-valued, are capable of capturing asymmetric distributions with heavy tails, as well as a wide range of serial correlation patterns, including both short and long memory behavior. A key advantage is that their statistical properties are fully characterized by their marginal distribution and autocorrelation function, both of which are available in closed form. The main difficulty, however, lies with parameter inference. Consider a trawl process $X$ parameterized by $\bm{\theta}\in\mathbb{R}^{m}$, sampled at discrete time points $1,\ldots,k$, yielding the observations $\bm{x}=(x_{1},\ldots,x_{k})$. Since trawl processes are generally not Markovian, and since the density $p(\bm{x}\mid\bm{\theta})$ is given by an intractable integral over k(k1)/2k(k-1)/2 dimensions (Leonte & Veraart 2023), classical maximum likelihood estimation is not feasible. For integer-valued trawls, Barndorff-Nielsen et al. (2014) proposed using generalized method of moments (GMM) estimators by matching theoretical and empirical moments and autocorrelations, while Bennedsen et al. (2023) and Leonte & Veraart (2023) developed pairwise likelihood (PL) estimators based on bivariate densities at fixed time lags. While PL estimators demonstrated superior finite-sample performance compared to GMM estimators in simulation studies, they face several limitations: unclear asymptotics in the long-memory case, unresolved optimal lag weighting, and computationally expensive Monte Carlo estimation of the bivariate densities. Furthermore, the PL approach cannot be easily extended beyond the univariate, stationary case. Given that trawl processes can be simulated efficiently using algorithms from Leonte & Veraart (2024), these challenges motivate the development of scalable simulation-based inference (SBI) methods that can compete with full likelihood-based estimation.
While SBI has traditionally focused on variants of approximate Bayesian computation (ABC), modern approaches have been dominated by the emergence of neural inference techniques. Given an observation $\bm{x}_{\text{o}}$, ABC generates many pairs $(\bm{x},\bm{\theta})$ and only retains these $\bm{\theta}$ for which $\bm{x}$ closely matches $\bm{x}_{\text{o}}$, i.e., $d\left(\bm{x},\bm{x}_{\text{o}}\right)<\epsilon$ for some metric $d(\cdot,\cdot)$ and error tolerance $\epsilon>0$. Several works attempt to address the impractically low ABC acceptance rates, by comparing informative summary statistics instead of the raw data (Blum et al. 2013) or by an MCMC-ABC extension (Marjoram et al. 2003), but the methodology remains sensitive to the choice of metric $d$ and tolerance $\epsilon$ (Robert et al. 2011), and is generally computationally inefficient. By contrast, neural inference uses simulations $(\bm{x},\bm{\theta})$ to train a neural network that directly approximates either (1) the likelihood $p\left(\bm{x}\mid\bm{\theta}\right)$, exactly (Papamakarios et al. 2017, 2019) or up to a constant (Hinton 2002), or (2) the full posterior $p\left(\bm{\theta}\mid\bm{x}\right)$ (Radev et al. 2022), or (3) point summaries thereof (Sainsbury-Dale et al. 2024, 2025, Richards et al. 2024). Once trained, the network can be evaluated on new observations without retraining, hence amortizing the initial training cost; see Zammit-Mangion et al. (2025) for a review of amortized inference with neural methods. Out of the available neural inference methodologies, we employ neural ratio estimation (NRE) because it avoids the complex architectures required by flow-based models and enables us to leverage binary classification tools to assess and improve the quality of the learnt approximation. Indeed, the main advantage of NRE is that it transforms the intractable density approximation task into a relatively simple binary classification problem, whereby a neural network learns to distinguish between samples from two carefully constructed distributions. The first is the joint distribution $p(\bm{x},\bm{\theta})=p\left(\bm{x}\mid\bm{\theta}\right)p(\bm{\theta})$, where $p(\bm{\theta})$ is a sampling distribution (often uniform) chosen to ensure good parameter space coverage during training. The second is the product of marginals $p(\bm{x})p(\bm{\theta})$, where $p(\bm{x})=\int p\left(\bm{x}\mid\bm{\theta}\right)p(\bm{\theta})\mathrm{d}\bm{\theta}$ is the marginal data distribution induced by the chosen sampling procedure. Crucially, $p(\bm{\theta})$ serves purely as a sampling distribution for generating training data, not as a prior distribution, and the actual prior used in Bayesian inference can be chosen independently. The trained classifier approximates the ratio $r(\bm{x},\bm{\theta})=p(\bm{x},\bm{\theta})/\left(p(\bm{x})p(\bm{\theta})\right)=p(\bm{x}\mid\bm{\theta})/p(\bm{x})$, which is proportional to the full likelihood for fixed $\bm{x}$, enabling both frequentist and Bayesian inference.
However, directly applying NRE to complex processes with stylized features, such as trawls, or to models with high-dimensional parameters $\bm{\theta}$ is challenging. In such cases, the joint density $p(\bm{x},\bm{\theta})$ is often highly concentrated relative to the product of marginals $p(\bm{x})p(\bm{\theta})$. Consequently, the ratio $r(\bm{x},\bm{\theta})$ takes extremely large values in a narrow region of the parameter space and is nearly zero elsewhere, and the classifier struggles to learn meaningful level sets. While it is well known that NRE can produce posterior approximations with poorly calibrated credible regions (Hermans et al. 2022), we uncover an additional critical failure: classical NRE also fails to accurately locate the likelihood mode itself. This dual failure, both in uncertainty quantification and point estimation represents a fundamental breakdown. We make four key contributions that address these challenges and advance SBI methodology along several fronts: (i) improved sample efficiency during training, (ii) reduced computational cost during inference, (iii) novel diagnostic tools to detect errors in the learnt likelihood, and (iv) post-calibration techniques that not only correct these errors but also enable further amortization over the sequence length $k$ of $\bm{x}$. Importantly, while we illustrate our approach in the context of trawl processes, it is generally applicable to any stochastic process for which standard SBI performs inadequately.
In our first contribution (i), we extend telescoping ratio estimation (TRE; Rhodes et al. 2020) to the SBI setting by decomposing the likelihood ratio as a product of $m$ simpler, one-dimensional conditional ratios $r(\bm{x},\bm{\theta})=\prod_{i=1}^{m}r_{i}\left(\bm{x},\bm{\theta}^{1:i}\right)$, where the term $r_{i}$ only depends on $\bm{\theta}\in\mathbb{R}^{m}$ through its first $i$ components, denoted $\bm{\theta}^{1:i}$. This decomposition simplifies the learning task, dramatically improving training efficiency. We indeed formally show that this setup requires exponentially fewer samples $(\bm{x},\bm{\theta})$ to reach the same approximation quality as standard NRE. In our second contribution (ii), we introduce a novel, MCMC-free, GPU-friendly sequential sampling approach that builds upon the proposed TRE decomposition and significantly accelerates posterior inference compared to state-of-the-art schemes such as the No U-Turn Sampler (NUTS). By accurately approximating each of the one-dimensional conditional densities $\left\{p\left(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1}\right)\right\}_{i=1}^{m}$ using Chebyshev polynomials (Olver & Townsend 2013), which converge uniformly on compacts, we enable efficient inverse transform sampling with a computational time that is linear in the dimensionality $m$ of the parameter $\bm{\theta}$. We also generalize this sequential approach to classifiers that model two components of $\bm{\theta}$ jointly. In our third contribution (iii), we introduce component-wise diagnostics that extend existing SBI validation methods and make them computationally efficient using the aforementioned Chebyshev polynomials. While this adds a second approximation layer on top of the TRE, which already approximates the true density, it crucially decouples expensive neural network evaluations from tasks such as posterior sampling or computation of highest posterior density regions. Finally, in our fourth contribution (iv), we show that even after breaking the learning task into easier ones through TRE, the classifiers still benefit from post-training calibration, as calibration is a global property and cannot be reliably enforced from batches during stochastic gradient descent. We also show that calibration can be further exploited to amortize the SBI estimator across different sequence lengths.
We illustrate the practical utility of our methodology and the benefits of amortization in a simulation study on trawl processes and an application to energy demand data.
Paper outline: Section 2 reviews trawl processes, and presents their key properties. Section 3.1 recalls background on the NRE framework for learning likelihood ratios, and discusses its main limitations. Section 3.2 introduces our novel SBI methodology based on the TRE decomposition and its theoretical foundations, while Section 3.3 presents the sequential inference strategy using Chebyshev polynomials. The remainder of Section 3 covers post-training calibration and diagnostics to assess the quality of learnt approximations. Sections 4 and 5 demonstrate the effectiveness of our approach through a simulation study on trawl processes and an application to real-world energy demand data. Section 6 concludes with an overview of implications for the broader SBI literature and directions for future research. Additional details and results are available in the Supplementary Material.
2 Modelling with Lévy bases and trawl processes
Let $\mathcal{B}_{\text{Leb}}(\mathbb{R}^{d})$ be the Borel $\sigma$-algebra on $\mathbb{R}^{d}$ restricted to sets of bounded Lebesgue measure.
2.1 Lévy bases
Definition 2.1.
A Lévy basis $L$ is a collection of infinitely-divisible, real-valued random variables $\left\{L(A):A\in\mathcal{B}_{\text{Leb}}(\mathbb{R}^{d})\right\}$ such that for any disjoint sets $A,B\in\mathcal{B}_{\text{Leb}}(\mathbb{R}^{d}),$ the random variables $L(A)$ and $L(B)$ are independent and $L\left(A\cup B\right)=L(A)+L(B)$ almost surely.
We focus on stationary Lévy bases, defined by the property that for any $\mathbf{z}\in\mathbb{R}^{d}$ and $A\subset\mathbb{R}^{d}$, $L(A+\mathbf{z})$ and $L(A)$ are equal in law, where $A+\mathbf{z}$ is the set addition. Barndorff-Nielsen et al. (2018) shows that stationary Lévy bases are homogeneous, meaning the law of $L(A)$ depends on $A$ only through its area $\textrm{Leb}(A)$. Further, Pedersen (2003) extends the Lévy-Ito decomposition from Lévy processes to bases, showing that $L$ can be written as the sum of independent Gaussian and jump Lévy bases as $L=L_{g}+L_{j}$, with $L_{g}(A)\sim\mathcal{N}\left(\mu\,\textrm{Leb}(A),\sigma^{2}\,\textrm{Leb}(A)\right)$ for some drift $\mu$, variance $\sigma^{2}$ and compensated jump process $L_{j}(A)$.
For convenience, we call a random variable $L^{\prime}$ a Lévy seed for $L$ if $L^{\prime}$ has the same law as $L\left([0,1]^{d}\right)$. More generally, any $L(A_{1})$ is again a Lévy seed for any set $A_{1}\in\mathcal{B}_{\text{Leb}}(\mathbb{R}^{d})$ with $\mathrm{Leb}(A_{1})=1$. The notion of Lévy seeds allows us to state the distributional properties of $L$ without reference to a specific set and provides a flexible class of marginal distributions, supported on the integers, reals or positive reals. A popular example encompassing many families of Lévy seeds is the class of generalized hyperbolic distributions (Borak et al. 2011), denoted $\textrm{GH}(\lambda,\alpha,\beta,\delta,\mu)$, which has semi-heavy tails and density given by
$x\mapsto\frac{\left(\gamma/\delta\right)^{\lambda}}{\sqrt{2\pi}K_{\lambda}(\delta\gamma)}\frac{K_{\lambda-1/2}\left(\alpha\sqrt{\delta^{2}+(x-\mu)^{2}}\right)}{\left(\sqrt{\delta^{2}+(x-\mu)^{2}}/\alpha\right)^{1/2-\lambda}}e^{\beta(x-\mu)},\text{ for }\alpha,\delta\in\mathbb{R}^{+}\text{ and }\beta,\mu,\lambda\in\mathbb{R},$
and where $\gamma\vcentcolon=\sqrt{\alpha^{2}-\beta^{2}}$ and $K_{\lambda}$ is the modified Bessel function of the second kind of order $\lambda$. This family includes many standard distributions that are supported on the reals as special or limiting cases and offers great flexibility. From a statistical modelling point of view, it is often advantageous to have the marginal law of $L(A)$ in closed form. Many common infinitely divisible distributions (e.g., Poisson, Gaussian, and Gamma) are closed under convolutions, meaning that $L^{\prime}$ and $L(A)$ are from the same named family: for example, if $L^{\prime}\sim\textrm{Poisson}(\nu)$, then $L(A)\sim\textrm{Poisson}(\nu\,\textrm{Leb}(A))$. While the GH family is not convolution-closed in general, it contains two subfamilies that are: the Normal-inverse Gaussian (NIG) distribution, corresponding to $\lambda=-1/2$, and the Variance Gamma (VG) distribution, for $\delta=0$. Indeed, if $L^{\prime}\sim\textrm{NIG}(\alpha,\beta,\delta,\mu)$, then $L(A)\sim\textrm{NIG}\left(\alpha,\beta,\delta\,\mathrm{Leb}\left(A\right),\mu\,\mathrm{Leb}\left(A\right)\right)$ and if $L^{{}^{\prime}}\sim\textrm{VG}(\alpha,\beta,\lambda,\mu)$, then $L(A)\sim\textrm{VG}\left(\alpha,\beta,\lambda\,\mathrm{Leb}(A),\mu\,\mathrm{Leb}(A)\right)$. These classes preserve closed-form marginals while allowing for both exponential and light tails. Parameterizations of all mentioned distributions, along with additional heavy-tailed examples and details on their scaling with $\mathrm{Leb}(A)$, are provided in Section LABEL:suppl_material:parameterizations.
2.2 Kernel convolutions and trawl processes
A general method for constructing stochastic processes $X$ driven by Lévy bases is based on kernel convolutions. Under mild regularity conditions on the kernel $K$, the convolution
$X(\cdot)=\int_{\mathbb{R}^{q}}K(\cdot,\bm{y})L(\mathrm{d}\bm{y}),\qquad q\in\mathbb{N},$
is well-defined (Rajput & Rosiński 1989, Theorem 2.7). To demonstrate our methodology, we focus on the univariate, time-only case. However, extensions to multivariate, spatio-temporal processes are straightforward (Barndorff-Nielsen et al. 2018, Opitz 2017). Of particular interest are indicator kernel functions, yielding the class of univariate trawl processes, originally introduced by Wolpert & Taqqu (2005) and Barndorff-Nielsen et al. (2014). Trawl processes are infinitely divisible, and their marginals often belong to the same family as the Lévy seed. Formally, taking $q=1$, consider the collection of sets
$A_{t}=A+(t,0),\qquad A=\{(s,y)\in\mathbb{R}^{2}\colon s<0,0<y<a(s)\},$
where $a\colon(-\infty,0]\to\mathbb{R}^{+}$ is a smooth, increasing function, and let the kernels be indicators over the sets $A_{t}$, i.e., $K(t,y)=\mathds{1}\left({(t,y)\in A_{t}}\right)$. Then the trawl process $X$ at time $t$, denoted by $X_{t}\equiv X(t)$, is given by the Lévy basis $L$ evaluated over the trawl set $A_{t}$, i.e., $X_{t}=L(A_{t})$. A key property of trawl processes is that their statistical properties are fully determined by the autocorrelation structure and marginal distribution. Therefore, trawl processes can be viewed as the natural extension of Gaussian processes to the infinitely divisible setting.
The flexibility of the marginal law of the trawl process is inherited from the Lévy basis. To maintain a flexible autocorrelation structure, the trawl sets incorporate an additional, abstract dimension $y$. By smoothing the Lévy basis in higher dimensions, trawl processes can achieve any smooth, decreasing autocorrelation function (Barndorff-Nielsen et al. 2014). Precisely, their correlation structure is given by
$\rho(h)\vcentcolon=\operatorname{Corr}(X_{t},X_{t+h})=\frac{\mathrm{Leb}\left(A\cap A_{h}\right)}{\mathrm{Leb}\left(A\right)}=\frac{\int_{-h}^{0}a(s)\mathrm{d}s}{\int_{-\infty}^{0}a(s)\mathrm{d}s},\text{ for }h\geq 0.$
The simplest parametric example is the exponential trawl function, where $a(s)=e^{\lambda s}$ for $s\leq 0,\lambda>0$, which yields $\rho(h)=e^{-\lambda h}$ for $h\geq 0$. To interpolate between short and long memory, consider the randomized mixture of exponentials $a(s)=\int_{0}^{\infty}e^{-\lambda s}\pi(\mathrm{d}\lambda)$ for $s\leq 0$, where the decay rate $\lambda$ is sampled according to $\pi$, a probability measure on $\mathbb{R}^{+}$. This recovers the purely exponential case with the Dirac measure $\pi=\delta_{\lambda_{0}}$ for some $\lambda_{0}>0$, and it also includes mixtures of exponentials for $\pi=\sum_{i=1}^{n}\lambda_{i}\delta_{\lambda_{i}}$, with $\sum_{i=1}^{n}\lambda_{i}=1$. In the simulations in Section 4, we use the Inverse Gaussian (IG) trawl function, where $\pi=\textrm{IG}(\gamma,\eta)$ for $\gamma,\,\eta>0$ and $a(s)=\left(1-2s/\gamma^{2}\right)^{-1/2}\exp{\left(\eta\left(1-\sqrt{1-2s/\gamma^{2}}\right)\right)}$, $s\leq 0$, yielding semi-long memory with $\rho(h)=\exp{\left(\eta\left(1-\sqrt{1+2h/\gamma^{2}}\right)\right)}$, $h\geq 0$. By semi-long memory we mean $\rho(\cdot)$ decays slower than exponential but faster than polynomial; see Section LABEL:suppl_material:parameterizations for parameterizations.
In summary, trawl processes are integer or real-valued, continuous-time stochastic processes that can describe a wide range of possible serial correlation patterns in data and whose properties are entirely determined by their marginal distribution and autocorrelation function. As a result of the latter, trawl processes and their spatio-temporal and multivariate extensions can be viewed as generalizations of Gaussian processes. Despite their appealing theoretical properties, parameter inference is challenging, as explained in the introduction. To this end, we leverage the fast simulation algorithms from Leonte & Veraart (2024) to enable simulation-based inference (SBI) techniques, specifically neural ratio estimation (NRE).
3 Simulation-based inference methodology
In this section, we first recall background on the NRE framework for SBI, which consists in learning likelihood ratios through classification, and discuss its main limitations. We then introduce our novel SBI methodology that combines TRE for improved training efficiency and a fast MCMC-free Bayesian inference approach based on Chebyshev polynomials. We also discuss post-calibration strategies to remedy the approximation error and enable amortization over the sequence length in the time series context. Finally, we propose novel tests and diagnostics to assess the quality of the learnt likelihood.
3.1 NRE: learning likelihood ratios through classification
[…]
$c^{*}\vcentcolon=\operatorname*{arg\,max}_{c}\,\mathbb{E}\left[\log{c(\mathbf{Z})}+\log{(1-c(\tilde{\mathbf{Z}}))}\right$]. (3.1)
By a calculus of variations argument (Cranmer et al. 2015, Proposition 2), it follows that
$c^{*}(\mathbf{z})=\frac{q(\mathbf{z})}{q(\mathbf{z})+\tilde{q}(\mathbf{z})}.$ (3.2)
Assuming the Bayes optimal classifier is available, the likelihood ratio can be recovered as
$r(\mathbf{z})\vcentcolon=\frac{q(\mathbf{z})}{\tilde{q}(\mathbf{z})}=\frac{q(\mathbf{z})/\left({q(\mathbf{z})+\tilde{q}(\mathbf{z})}\right)}{\tilde{q}(\mathbf{z})/\left({q(\mathbf{z})+\tilde{q}(\mathbf{z})}\right)}=\frac{c^{*}(\mathbf{z})}{1-c^{*}(\mathbf{z})}.$
In practice, neither the BCE loss from (3.1) nor $c^{*}$ are available. In lieu, consider a flexible set of parametric classifiers $\{c_{\bm{\psi}}:\bm{\psi}\in\Psi\}$ satisfying a universal approximation theorem, e.g., based neural networks, and minimize the Monte Carlo approximation of the BCE loss
$\frac{1}{2N}\left[\sum_{i=1}^{N}\log{\left(c_{\bm{\psi}}(\mathbf{z}_{i})\right)}+\sum_{i=1}^{N}\log{\left(1-c_{\bm{\psi}}(\tilde{\mathbf{z}}_{i})\right)}\right],$ (3.3)
where $\left\{\mathbf{z}_{i}\right\}_{i=1}^{N}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}q$ and $\left\{\tilde{\mathbf{z}}_{i}\right\}_{i=1}^{N}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\tilde{q}$, to obtain a minimizer $\hat{\bm{\psi}}$ and the corresponding approximations $\hat{c}\vcentcolon=c_{\hat{\bm{\psi}}}$ of $c^{*}$ and $\hat{r}$ of $r$. Cranmer et al. (2015) and Hermans et al. (2020) pioneer the above classification-based approach within SBI, when fast samplers are available. Specifically, let $\bm{x}\in\mathcal{X}$ be a realization of a stochastic process parameterized by $\bm{\theta}\in\Theta$—in our case a trawl process—and let $\mathbf{z}=(\bm{x},\bm{\theta})\in\mathcal{Z}=\mathcal{X}\times\Theta\subset\mathbb{R}^{k+m}$. For ease of notation, we write $p(\bm{x}\mid\bm{\theta})$ for the likelihood function in both the frequentist and Bayesian case, and explain the differences below. Consider a sampling density $p(\bm{\theta})$ on $\Theta$ and the induced joint, marginal, and product of marginal densities $q(\bm{x},\bm{\theta})\vcentcolon=p(\bm{x},\bm{\theta})=p(\bm{x}\mid\bm{\theta})p(\bm{\theta})$, $p(\bm{x})=\int p(\bm{x},\bm{\theta})\mathrm{d}\bm{\theta}$, and $\tilde{q}(\bm{x},\bm{\theta})\vcentcolon=p(\bm{x},\bm{\theta})=p(\bm{x})p(\bm{\theta})$, respectively. A classifier $\hat{c}$ trained to distinguish between samples from $q$ and $\tilde{q}$ yields an approximation $(\bm{x},\bm{\theta})\mapsto\hat{r}(\bm{x},\bm{\theta})$ of
$r(\bm{x},\bm{\theta})=\frac{q(\bm{x},\bm{\theta})}{\tilde{q}(\bm{x},\bm{\theta})}=\frac{p(\bm{x},\bm{\theta})}{p(\bm{x})p(\bm{\theta})}=\frac{p(\bm{x}\mid\bm{\theta})}{p(\bm{x})}\ \propto\ p(\bm{x}\mid\bm{\theta})\text{ for fixed }\bm{x}.$
Thus, NRE facilitates both frequentist and Bayesian inference. In the former, $\hat{r}$ approximates the likelihood up to a normalizing constant. In the latter, the posterior is approximated via Bayes’ rule as $p(\bm{\theta}\mid\ \bm{x})=r(\bm{x},\bm{\theta})p(\bm{\theta})\approx\hat{r}(\bm{x},\bm{\theta})p(\bm{\theta})$ if $p(\bm{\theta})$ is also used as the prior. The prior used for inference does not have to match the sampling distribution used during training. In this case, the posterior is available up to a constant. Importantly, the approximation $r\approx\hat{r}$ is valid only in regions of the parameter space well covered by training samples. For this reason, the sampling density $p(\bm{\theta})$ is often chosen as a product of uniform distributions.
[…]
3.2 TRE: a novel divide-and-conquer SBI technique
To address the above issues, we build upon on Rhodes et al. (2020) and extend telescoping ratio estimation (TRE) to the SBI context, showing that this novel approach requires exponentially fewer samples compared to NRE. The main idea is to introduce suitable interpolating distributions $q_{0},\ldots,q_{m}$, such that $q_{0}=\tilde{q}$ and $q_{m}=q$, and to learn the likelihood-ratio by training $m$ classifiers and multiplying the corresponding ratios. Writing $\bm{\theta}=(\theta^{1},\ldots,\theta^{m})\in\Theta\subset\mathbb{R}^{m}$ and $\bm{\theta}^{i:j}=(\theta^{i},\ldots,\theta^{j})$ for any integers $1\leq i\leq j\leq m$, define
$q_{i}(\bm{x},\bm{\theta})=p(\bm{x},\bm{\theta}^{1:i})p(\bm{\theta}^{i+1:m}),\qquad i=1,\ldots,m-1,$
and $q_{m}(\bm{x},\bm{\theta})=p(\bm{x},\bm{\theta})$ and $q_{0}(\bm{x},\bm{\theta})=p(\bm{x})p(\bm{\theta})$. Furthermore, define the density ratios
$r_{i}(\bm{x},\bm{\theta})\vcentcolon=\frac{q_{i}(\bm{x},\bm{\theta})}{q_{i-1}(\bm{x},\bm{\theta})}=\frac{p(\bm{x},\bm{\theta}^{1:i})p(\bm{\theta}^{i+1:m})}{p(\bm{x},\bm{\theta}^{1:i-1})p(\bm{\theta}^{i:m})}=\frac{p(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})}{p(\theta^{i}\mid\bm{\theta}^{i+1:m})},\qquad i=1,\ldots,m,$
with the convention that $\bm{\theta}^{1:0}=\bm{\theta}^{m+1:m}=\emptyset$. Note that the desired ratio $r(\bm{x},\bm{\theta})=p(\bm{x},\bm{\theta})/\left(p(\bm{x})p(\bm{\theta})\right)$ can then be obtained as $r(\bm{x},\bm{\theta})=\prod_{i=1}^{m}r_{i}(\bm{x},\bm{\theta})$. The main benefit is that the density ratios $r_{i}$ can be approximated separately, by training $m$ classifiers to distinguish between samples from $q_{i}$ and $q_{i-1}$, for $i=1,\ldots,m$. We propose the architecture in Figure 1(a) for each of the $m$ classifiers. The choice of the interpolating distributions $\{q_{i}\}_{i=0}^{m}$ is directly linked to training efficiency, as revealed by Theorem 3.1.
(a) Architecture of $i^{\text{th}}$ classifier within TRE.
(b) TRE architecture with shared encoder.
Figure 1: TRE architectures for learning likelihood ratios. Outputs are on the log scale for stability. (a) Independent classifiers: Each encoder (e.g., Long Short-Term Memory, or LSTM) processes the data (e.g., time series) $\bm{x}$ into summary statistics $s_{i}(\bm{x})$, which are concatenated with $\bm{\theta}$ and passed to a multilayer perceptron (MLP) to approximate $\log{\hat{p}(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})}$. This is added to $\log{p(\theta^{i}\mid\bm{\theta}^{i+1:m})}$ to yield $\log\hat{r}_{i}(\bm{x},\bm{\theta})$. (b) Shared encoder variant: All classifiers share a single encoder with separate MLP heads; see Section LABEL:SM:comparison_with_original_TRE. The $\oplus$ symbol indicates addition by $\log p(\theta^{i}|\bm{\theta}^{i+1:m})$.
Theorem 3.1 (Sample efficiency).
Let $q_{0}(\bm{x},\bm{\theta})=p(\bm{x})p(\bm{\theta})$, $q_{m}(\bm{x},\bm{\theta})=p(\bm{x},\bm{\theta})$ and $q_{i}(\bm{x},\bm{\theta})=p(\bm{x},\bm{\theta}^{1:i})p(\bm{\theta}^{i+1:m})$ for $i=1,\ldots,m-1$, as above. If $p(\bm{\theta})=p(\theta^{1})\cdots p(\theta^{m})$, then
$D_{\textrm{KL}}(p(\bm{x},\bm{\theta})\;\|\;p(\bm{x})p(\bm{\theta}))=D_{\textrm{KL}}(q_{m}\;\|\;q_{0})=\sum_{i=1}^{m}D_{\textrm{KL}}(q_{i}\;\|\;q_{i-1}).$
See Section LABEL:appendix:subsection_tre for the proof and the general case where $p(\bm{\theta})$ does not factorize. While NRE requires exponentially many samples in $D_{\textrm{KL}}(q_{m}\;\|\;q_{0})$ to accurately represent the BCE loss (Chatterjee & Diaconis 2018), the $i^{\text{th}}$ TRE classifier only requires exponentially many samples in $\mathcal{S}_{i}\vcentcolon=D_{\textrm{KL}}(q_{i}\;\|\;q_{i-1})$. Since all $m$ classifiers can be trained using the same simulated $(\bm{x},\bm{\theta})$ samples, the overall sample complexity scales exponentially in $\max_{1\leq i\leq m}\mathcal{S}_{i}$. Crucially, the order of parameter estimation matters. We assign $\theta^{1}$ to the most influential component and $\theta^{m}$ to the least, since learning $\theta^{m}$ is conditioned on $\bm{\theta}^{1:i-1}$ and is therefore easier, while $\theta^{1}$ must be learned in isolation. This ordering roughly brings the $\mathcal{S}_{i}$ values to similar magnitudes and approximately yields an effective $1/m$ reduction in the exponent. If large discrepancies in $\mathcal{S}_{i}$ appear during training, reordering accordingly improves performance. Note that the terms $\mathcal{S}_{i}=D_{\textrm{KL}}(q_{i}\;\|\;q_{i-1})=\mathbb{E}_{q_{i}(\bm{x},\bm{\theta})}\left[\log{r}_{i}(\bm{x},\bm{\theta})\right]$ can be estimated during training directly from the classifier outputs, with no additional computation. Finally, the $\mathcal{S}_{i}$ values can also inform the allocation of batch sizes across individual classifiers.
Remark 3.2.
So far, we treated each scalar parameter $\theta^{i}$ in isolation, by learning one-dimensional conditional densities $p(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})$. In practice, however, it may be advantageous to group into blocks these components of $\bm{\theta}$ that are interlinked and cannot be made orthogonal. Concretely, we can partition $\bm{\theta}\in\mathbb{R}^{m}$ into $l<m$ subvectors as $\bm{\theta}=(\bm{\theta}^{(1)},\dots,\bm{\theta}^{(l)})$, and learn the block‐conditional densities $p(\bm{\theta}^{(j)}\mid\bm{x},\bm{\theta}^{(1:j-1)})$, $j=1,\ldots,l$. This strategy is useful when some components of $\bm{\theta}$ are easier to identify jointly rather than separately, and may improve the quality of the ratio estimates. Theorem 3.1 stays true with a similar proof.
Remark 3.3.
We stress that our approach differs fundamentally from Rhodes et al. (2020). While the original TRE therein targets direct density estimation by learning a neural approximation $\hat{p}_{\bm{\psi}}(\cdot)$ of $\bm{x}\mapsto p(\bm{x})$ for a given dataset without an underlying parametric model, our approach is embedded within the SBI framework and learns an amortized approximation $\hat{p}_{\bm{\psi}}(\cdot\mid\cdot)$ of $(\bm{x},\bm{\theta})\mapsto p(\bm{\theta}\mid\bm{x})$. More importantly, the original formulation suffers from a training-inference mismatch due to ratio evaluations in regions of negligible density under the training data, which our approach remedies; see Section LABEL:SM:comparison_with_original_TRE for details.
3.3 MCMC-free posterior sampling via Chebyshev polynomials
Fast posterior sampling is crucial for both practical applications and the coverage checks from Section 3.5, which require generating posterior samples for thousands of model realizations. Obtaining satisfactory effective sample sizes in this setting, in particular for the simulation study in Section 4, requires adaptive samplers such as NUTS, which tune not only the proposal covariance and step size, but also the number of leapfrog steps. Even this scheme proves inefficient: the adaptive nature prevents any GPU acceleration, makes each chain computationally costly, and forces chains to be run sequentially. As a result, carrying out coverage checks with MCMC is impractical, motivating the development of a novel, MCMC-free approach. However, it is worth noting that when gradient-based MCMC is viable—or when performing maximum likelihood estimation via gradient descent—we can exploit the TRE framework to halve the number of gradient evaluations; see Section LABEL:SM:reduced_gradient_cost.
Having gained access to the one-dimensional conditional densities $\left\{p(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})\right\}_{i=1}^{m}$, we can generate posterior samples sequentially in the components of $\bm{\theta}$. Instead of MCMC, we propose an efficient inversion sampling algorithm, by constructing approximate cumulative distribution functions (CDFs) $\hat{F}\left(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1}\right)$ which are simple to evaluate and then invert by bisection. A detailed account is provided in Section LABEL:suppl:subsection_chebyshev, and we only summarize the main idea here. Following Olver & Townsend (2013), we approximate the one-dimensional densities using Chebyshev polynomials, which only require evaluations at specific Chebyshev knots. This approach has several appealing properties. Computationally, these evaluations only have to be carried out once, regardless of the required number of samples, can be efficiently parallelized on a GPU, do not involve a computationally expensive encoder, and perform reliably in practice. Theoretically, the Chebyshev polynomial approximations converge uniformly on compact intervals and allow both the polynomial and its integral (the approximate CDF) to be evaluated efficiently in closed form, via Clenshaw’s algorithm (Clenshaw 1955), without additional neural network evaluations. Finally, this approach allows us to introduce novel, computationally efficient checks to individually test each component classifier, as explained in Section 3.5. Because the generated samples are independent, we avoid thinning procedures that are often necessary with MCMC samples for simulation-based calibration diagnostics such as rank-based checks (Talts et al. 2018).
3.4 Calibration analysis and further amortization
3.4.1 Potential reasons for miscalibration and remedies
We say that a binary classifier $c:\mathcal{Z}\to[0,1]$ is calibrated if the following holds
$\mathds{P}\left(Y=1\mid c(\mathbf{Z})\right)=c(\mathbf{Z})\ \text{ a.s. over pairs }(\mathbf{Z},Y),$
where $Y\in\{0,1\}$ denotes the true label of data $\mathbf{Z}$. Intuitively, among all inputs $\mathbf{z}$ for which $c(\mathbf{z})=0.75$, say, about $75\%$ should have label $Y=1$. Nevertheless, neural network-based classifiers, including NRE-based ones, are often miscalibrated (Guo et al. 2017, Hermans et al. 2022). Miscalibration is typically quantified using the expected calibration error (ECE), $\mathbb{E}\left[\left|\mathds{P}\left(Y=1\mid c(\mathbf{Z})\right)-c(\mathbf{Z})\right|\right]$, which measures the average discrepancy between the predicted confidence $c(\mathbf{Z})$ and the accuracy $\mathds{P}\left(Y=1\mid c(\mathbf{Z})\right)$, and can be approximated from samples as discussed in Section LABEL:SM:ECE. Empirical studies (Niculescu-Mizil & Caruana 2005, Guo et al. 2017, Minderer et al. 2021) report conflicting results, with calibration varying based on regularization, architecture and model capacity. We identify two main avenues for improvement and further leverage the rich literature on post-hoc calibration of probabilistic classifiers to ensure that the level curves of the likelihood (posterior) are calibrated.
The first issue is overparameterization. When the BCE loss plateaus, networks often increase confidence on correctly classified samples and output extreme values, near 0 or 1 (Mukhoti et al. 2020). The overfit occurs because the finite-sample BCE approximation is minimized when the network perfectly separates samples with extreme outputs, unlike the theoretical BCE loss, which is minimized by the optimal decision function from (3.2). This is formally analyzed by Soudry et al. (2018) for linearly separable data and is particularly problematic for time series, where longer time series enable near-perfect separation. We address this by leveraging the algorithms from Leonte & Veraart (2024) to simulate training data on-the-fly.
The second issue is intrinsic to classification itself. Bai et al. (2021) prove that logistic regression is overconfident even in ideal conditions, i.e., when the model is under-parameterized, the data follow the true logistic model, and the sample size far exceeds the parameter count. Although miscalibration vanishes asymptotically with dataset size, the computational cost of performing gradient descent iterations on a sufficiently large dataset is impractical. Therefore, post-training calibration is essential for obtaining reliable probability estimates.
3.4.2 Post-hoc calibration and amortization
Recall that $c(\bm{x},\bm{\theta})=p(\bm{x},\bm{\theta})/\left(p(\bm{x},\bm{\theta})+p(\bm{x})p(\bm{\theta})\right)$ and that classifier miscalibration directly propagates to the likelihood, posterior distribution, and credible region approximations, as $p(\bm{x}\mid\bm{\theta})\ \propto\ {p(\bm{\theta}\mid\bm{x})}\ \propto\ r(\bm{x},\bm{\theta})={\left[\left(c(\bm{x},\bm{\theta})\right)^{-1}-1\right]}^{-1}$ for fixed $\bm{x}$. To mitigate miscalibration, we apply a post-training monotonic transformation $T\colon[0,1]\to[0,1]$ to the trained classifier $\hat{c}$, yielding $\hat{c}_{\text{cal}}=T\circ\hat{c}$. The transformation is estimated using a separate dataset, allowing it to enforce global properties of the classifier that are difficult to capture from mini-batches during training. Based on Theorem $1$ from Cranmer et al. (2015), if $c$ is monotonic with respect to $c^{*}$, in the sense that $c(\bm{x},\bm{\theta}_{1})<c(\bm{x},\bm{\theta}_{2})$ implies $c^{*}(\bm{x},\bm{\theta}_{1})<c^{*}(\bm{x},\bm{\theta}_{2})$, then the optimal $c^{*}$ can be recovered from $c$ by calibration. In practice, we cannot guarantee the monotonicity condition, and we resort to the metrics and checks discussed in the next section to compare performance pre- and post-calibration.
Another advantage of calibration is amortization: in the time series context, once a classifier has been trained on a time series of length $k$, it can be reused on inputs with other lengths $k^{\prime}$ by fitting the calibration map on a newly simulated calibration dataset with $\bm{x}$ of length $k^{\prime}$. Intuitively, as $k$ increases, the classifier should be more confident, and the calibration map thus compensates by applying a monotonic correction, which naturally preserves the ranking of likelihoods, and Theorem $1$ from Cranmer et al. (2015) guarantees no loss in efficiency. In Section 4, we empirically verify this by showing that our TRE yields high-quality approximations even when applied to both shorter and longer sequences than those seen during training. The key is that calibration is fast and does not require retraining.
A popular parametric calibration method is Platt scaling (Platt et al. 1999), which applies the sigmoid transformation $T(s;A,B)=1/\left(1+\exp{(-As+B)}\right)$, with $A>0,B\in\mathbb{R}$. However, Platt scaling is quite rigid: it does not contain the identity, and can sometimes even worsen calibration. For this reason, Kull et al. (2017) introduce the more flexible family of beta-calibration mappings $T(s;a,b,c)=1/\left(1+e^{-c}(1-s)^{b}s^{-a}\right)$, with $a,b>0,c\in\mathbb{R}$, which yields the identity for $a=b=1$, $c=0$. Beyond parametric approaches, isotonic regression (Niculescu-Mizil & Caruana 2005) offers a more flexible, non-parametric alternative. It learns a monotone, stepwise function by solving a quadratic program and can yield superior performance. However, the resulting calibration mapping is piecewise constant, making it unsuitable when differentiability is required, e.g., in gradient-based MCMC. Fortunately, the MCMC-free posterior sampling scheme introduced for TRE in Section 3.3 lifts this requirement. We test both beta and isotonic calibration for comparison.
3.5 Checks on the quality of the approximate likelihood
3.5.1 Coverage diagnostics
Having trained and calibrated an amortized model, we next assess the quality of the approximation $(\bm{x},\bm{\theta})\mapsto\hat{p}(\bm{x}\mid\bm{\theta})$. Some approaches include classifier-based tests (Lopez-Paz & Oquab 2017, Linhart et al. 2023) and kernelized Stein divergence tests (Liu et al. 2016, Chwialkowski et al. 2016), though Lueckmann et al. (2021) found the above to be sensitive to hyperparameters and even inconsistent with each other. In this paper, we assess the approximation quality by comparing the theoretical and empirical coverage, i.e., we check whether ground-truth parameters fall within prediction regions at the correct rate.
Definition 3.4.
Let $\Theta_{\hat{p}(\cdot\,\mid\bm{x})}(1-\alpha)$ be the $1-\alpha$ highest posterior density region (HPD) of the approximate posterior $(\bm{x},\bm{\theta})\mapsto\hat{p}(\bm{\theta}\mid\bm{x})$. The expected coverage at level $1-\alpha$ is
$\mathcal{C}_{1-\alpha}=\mathbb{E}_{p(\bm{x},\bm{\theta})}\left[\mathds{1}\left({\bm{\theta}}\in\Theta_{\hat{p}(\cdot\,\mid\bm{x})}(1-\alpha)\right)\right].$ (3.4)
We stress that (3.4) can be thought of as both a frequentist expected coverage and a Bayesian credible region, depending on the interpretation of $\mathbb{E}_{p(\bm{x},\bm{\theta})}$ as $\mathbb{E}_{p(\bm{\theta})}\mathbb{E}_{p(\bm{x}\mid\bm{\theta})}$ or $\mathbb{E}_{p(\bm{x})}\mathbb{E}_{p(\bm{\theta}\mid\bm{x})}$. If the likelihood approximation $(\bm{x},\bm{\theta})\to\hat{p}(\bm{x}\mid\bm{\theta})$ is faithful, or equivalently if the posterior $(\bm{x},\bm{\theta})\to\hat{p}(\bm{\theta}\mid\bm{x})$ is well calibrated, then $\mathcal{C}_{1-\alpha}=1-\alpha$ for any $\alpha\in[0,1]$; discrepancies can be visualised by plotting the empirical and theoretical coverage against each other; see Figure 4 for illustration. The empirical coverage can be derived as follows. First, generate $N$ samples $(\bm{x}_{1},\bm{\theta}_{1}),\ldots,(\bm{x}_{N},\bm{\theta}_{N})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}p(\bm{x},\bm{\theta})$. For each $\bm{x}_{j}$, draw $M$ posterior samples $\bm{\vartheta}_{j,1},\ldots,\bm{\vartheta}_{j,M}\sim\hat{p}(\bm{\theta}\mid\bm{x}_{j})=\hat{r}(\bm{x}_{j}\mid\bm{\theta})\,p(\bm{\theta})$, e.g., using Chebyshev polynomials, and derive the approximate HPD region ${\Theta}^{M}_{\hat{p}(\cdot\,\mid\bm{x}_{j})}(1-\alpha)\approx{\Theta}_{\hat{p}(\cdot\,\mid\bm{x}_{j})}(1-\alpha)$ as follows: sort the values $p\left(\bm{\vartheta}_{j,1}\mid\bm{x}_{j}\right),\ldots,p\left(\bm{\vartheta}_{j,M}\mid\bm{x}_{j}\right)$ in descending order and take the top $(1-\alpha)M$ of them. The posterior density of the last included sample is then the threshold used to determine acceptance to the HPD. Finally, compare this threshold with $p(\bm{\theta}_{j}\mid\bm{x}_{j})$ and calculate the proportion of true parameters that fall within their corresponding HPD regions
$\hat{C}_{1-\alpha}=\sum_{j=1}^{N}\mathds{1}\left(\bm{\theta}_{j}\in\Theta^{M}_{\hat{p}(\cdot\,\mid\bm{x}_{j})}(1-\alpha)\right)=\sum_{j=1}^{N}\mathds{1}\left(p(\bm{\theta}_{j},\bm{x}_{j})\geq\textrm{threshold}_{j}\right).$ (3.5)
We emphasize that the $\Theta^{M}_{\hat{p}(\cdot\,\mid\bm{x}_{j})}(1-\alpha)$ region is doubly approximate: it is estimated via samples from the $\hat{p}(\cdot\mid\bm{x})$, which is itself an approximation of the posterior.
3.5.2 Posterior coverage limitations and novel individual checks
The coverage check may fail to detect poor approximations, particularly non-informative likelihoods. The degenerate case $\hat{p}(\bm{x}\mid\bm{\vartheta})\equiv 1$ results in $\hat{p}(\bm{\theta}\mid\bm{x})=p(\bm{\theta})$ and $\mathcal{C}_{1-\alpha}=1-\alpha$, but can be spotted by inspecting the classifier outputs, which are near $0.5$. More subtle failures evade immediate detection, such as when classifiers ignore specific components of $\bm{\theta}$ or when overconfidence and underconfidence across TRE components cancel out, yielding good coverage despite misspecified component-wise likelihoods.
To address these limitations, we introduce novel and computationally efficient per-parameter coverage checks. Our approach leverages the TRE framework’s access to the conditional densities $\hat{p}(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})$, which in turn enables the construction of HPD regions conditional on both $\bm{x}$ and preceding parameters $\bm{\theta}^{1:i-1}$, rather than $\bm{x}$ alone; once the HPD region is known, coverage is estimated as before, via Equation 3.5. We then estimate coverage across simulations as in (3.5). The key computational advantage lies in our Chebyshev polynomial approach: with typically fewer than $64$ neural network evaluations, we can accurately approximate $\hat{p}(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})$, generate arbitrary numbers of samples and compute HPDs instantaneously. Indeed, as opposed to MCMC, posterior samples do not require extra neural network evaluations once the coefficients of the Chebyshev polynomial are determined. Our experiments show sufficient accuracy, eliminating the need for importance sampling reweighting by the ratio between $\hat{p}(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})$ and its Chebyshev approximation. There, we also discuss the case where we infer blocks rather than scalar components of $\bm{\theta}$, as introduced in Remark 3.2. Overall, our component-wise approach enables targeted identification of miscalibrated classifiers, as well as model comparison.
3.5.3 Unified assessment: calibration, coverage, and metrics
As previously discussed, many SBI methodologies, including NRE, produce overly-peaky likelihoods, and equivalently overconfident posteriors (Hermans et al. 2022). In turn, this is equivalent to the overconfidence of the classifier, as $p(\bm{\theta}\mid\bm{x})\ \propto\ r(\bm{x},\bm{\theta})={\left[\left(c(\bm{x},\bm{\theta})\right)^{-1}-1\right]}^{-1}$. While post-hoc calibration can mitigate miscalibration, it can also degrade performance, e.g., for NREs distinguishing between samples which can be trivially separated. Hence, it is necessary to carry a case-by-case evaluation and selectively retrain when post-hoc calibration alone cannot achieve satisfactory diagnostic results. To aid in this analysis, we track and compare various metrics during training, as well as before and after calibration.
While no single metric fully characterizes posterior quality, monitoring several metrics during training helps identify issues and assess convergence. The BCE loss and classification accuracy directly measure classifier performance. The averaged cross-entropy $\mathcal{S}=\mathbb{E}_{p(\bm{\theta},\bm{x})}\left[\log{\hat{p}(\bm{\theta}\mid\bm{x})}\right]=\mathbb{E}_{p(\bm{x})}\mathbb{E}_{p(\bm{\theta}\mid\bm{x})}\left[\log{\hat{p}(\bm{\theta}\mid\bm{x})}\right]$ metric further enables comparison between SBI models, though its logarithmic nature makes it more sensitive to matching density peaks than level curves. Deviations of the balancing metric $\mathcal{B}=\mathbb{E}_{p(\bm{x},\bm{\theta})}\left[\hat{c}(\bm{x},\bm{\theta})\right]+\mathbb{E}_{p(\bm{x})p(\bm{\theta})}\left[\hat{c}(\bm{x},\bm{\theta})\right]$ from the theoretical value $1$ indirectly detects discrepancies of the aforementioned level curves through class imbalances. Finally, the ECE metric can serve for model comparison.
4 Simulation study
While we here focus on trawl processes, our methodology applies broadly to intractable stochastic processes that are otherwise difficult to tackle with vanilla SBI techniques.
4.1 Simulation setting
We apply our methodology to a class of trawl processes $X$ capable of exhibiting short and long memory, semi-heavy tails and skewness. To this end, consider $X_{t}\sim\textrm{NIG}(\mu,\sigma,\beta)$ with autocorrelation function $\rho(h)=\exp{\left(\,\eta_{\textrm{acf}}\left(1-\sqrt{1+2h/\gamma_{\textrm{acf}}^{2}}\right)\right)}$ for $h\geq 0$. We adopt a three-parameter form for the NIG distribution, see Section LABEL:subsubsection:alternative_specification, instead of the conventional four-parameter form for two reasons. First, under the standard four-parameter formulation, distinct parameter values can yield densities that are visually indistinguishable, hindering inference and posterior sampling. Second, our reparameterization remains interpretable, as $\mu$ and $\sigma$ correspond to the mean and standard deviation. We set independent uniform sampling distributions to ensure the amortized model is valid for a wide range of cases: $\gamma_{\text{acf}}\sim\mathcal{U}(10,20),\,\eta_{\text{acf}}\sim\mathcal{U}(10,20),\,\mu\sim\mathcal{U}(-1,1),\,\sigma\sim\mathcal{U}(0.5,1.5),\,\beta\sim\mathcal{U}(-5,5)$, and let $\bm{\theta}=\left(\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu,\sigma,\beta\right)$. The support restrictions on $\mu$ and $\sigma$ are without loss of generality, as we can feed the classifier time series that are centered and scaled with the empirical mean and standard deviation, and invert the transformation afterwards.
As discussed in Section 3.2, coordinate ordering matters in TRE. The autocorrelation functions (ACF) parameters are weakly identifiable, hence we infer these jointly. We thus learn a block of two coordinates, and we place it first in the TRE decomposition below to reduce the computational cost of building the Chebyshev approximation; see Section LABEL:SM:seq_p_sam. We then order the remaining parameters by expected learning difficulty: mean, standard deviation, then tilt. Formally, let $\bm{x}$ be an observation of $X$ at times $1,\ldots,k=1500$. Then, the ratio $r(\bm{x},\bm{\theta})$ can be learnt by training four classifiers:
$r(\bm{x},\bm{\theta})$ $=\frac{p(\bm{x},\bm{\theta})}{p(\bm{x})p(\bm{\theta})}=\frac{p(\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu,\sigma,\beta)}{p(\bm{x})\,p(\gamma_{\textrm{acf}},\eta_{\textrm{acf}})\,p(\mu)\,p(\sigma)\,p(\beta)}$
$=\underbrace{\frac{p(\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu,\sigma,\beta)}{p(\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu,\sigma)\,p(\beta)}}_{\beta\text{ classifier}}\cdot\underbrace{\frac{p(\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu,\sigma)}{p(\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu)\,p(\sigma)}}_{\sigma\text{ classifier}}\cdot\underbrace{\frac{p(\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu)}{p(\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}})\,p(\mu)}}_{\mu\text{ classifier}}\cdot\underbrace{\frac{p(\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}})}{p(\bm{x})\,p(\gamma_{\textrm{acf}},\eta_{\textrm{acf}})}}_{\text{ACF classifier}}$
$=\underbrace{\frac{p(\beta\mid\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu,\sigma)}{p(\beta)}}_{\beta\text{ classifier}}\cdot\underbrace{\frac{p(\sigma\mid\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu)}{p(\sigma)}}_{\sigma\text{ classifier}}\cdot\underbrace{\frac{p(\mu\mid\bm{x},\gamma_{\textrm{acf}},\eta_{\textrm{acf}})}{p(\mu)}}_{\mu\text{ classifier}}\cdot\underbrace{\frac{p(\gamma_{\textrm{acf}},\eta_{\textrm{acf}}\mid\bm{x})}{p(\gamma_{\textrm{acf}},\eta_{\textrm{acf}})}}_{\text{ACF classifier}}.$
We monitor the BCE loss, cross-entropy $\mathcal{S}$, accuracy, and balancing metric $\mathcal{B}$ (see Section 3.5.3) throughout training and display the results in Figure 2. All metrics approximately stabilize, suggesting convergence. Gradients are computed with data simulated on-the-fly at each training iteration, whereas the metrics are evaluated on a fixed, holdout dataset. We use the architecture from Figure 1(a), in which trawl process realizations $\bm{x}$ are fed through an LSTM encoder. The encoded features are concatenated with the relevant parameter subset of $\bm{\theta}$ and then passed through fully connected layers; implementation details are available in Section LABEL:section:training_details.
Figure 2: BCE, $\mathcal{S}$, accuracy, and $\mathcal{B}$ metrics (left to right) for the $\textrm{ACF},\,\mu,\,\sigma$ and $\beta$ classifiers (different colored lines), evaluated on a holdout dataset over the last 35000 training iterations. We train the classifiers with trawl process realizations $\bm{x}$ of length $1500$. The legend is displayed in the right panel.
Having obtained an approximate mapping $(\bm{x},\bm{\theta})\mapsto\hat{p}(\bm{x},\bm{\theta})$, we assess its quality when used for neural point estimation and for constructing posterior credible regions. As we show next, our methodology yields substantial improvements over existing approaches across three key dimensions: reduced estimation error, enhanced diagnostic capabilities for assessing the approximation’s quality, and amortization across varying sequence lengths.
4.2 Neural point estimators
We begin by numerically deriving maximum likelihood estimates (MLE) from the TRE approximation. For comparison, we consider two benchmarks: MLEs obtained from standard NRE, and point estimates from the generalized method of moments (GMM). For GMM, the marginal parameters are inferred from the first four empirical moments and ACF parameters from the empirical autocorrelation function. Since these methods serve only as benchmarks, we provide their implementation details in Section LABEL:sm:point_estimators. Figure 3 illustrates the performance of these methods on a trawl process realization of length $1500$. Note that TRE achieves the closest match to both the true marginal distribution and the ACF structure.
Figure 3: Performance comparison of point estimators given by TRE, NRE and GMM. Left: trawl process realization corresponding to $\bm{\theta}=(13.36,15.52,0.97,0.98,-0.17)$; middle: true (dashed) and inferred (solid) marginal distributions; right: true (dashed), empirical (solid-dotted), and infered (solid) ACFs.
    ACF $\mu$ $\sigma$ $\beta$ mean KL
    mean $L^{1}$ mean $L^{2}$    MAE RMSE    MAE RMSE    MAE RMSE
GMM     3.473 0.615    0.224 0.335    0.248 0.323    1.390 1.963    0.444
NRE     1.518 0.269    0.110 0.150    0.112 0.147    0.760 1.058    0.036
TRE     1.266 0.224    0.098 0.134    0.090 0.119    0.631 0.860    0.026
1000 NBE     1.218 0.215    0.101 0.135    0.089 0.115    0.529 0.695    0.040
GMM     3.084 0.546    0.196 0.300    0.226 0.299    1.285 1.834    0.374
NRE     1.308 0.232    0.094 0.125    0.098 0.127    0.686 0.964    0.025
TRE     1.071 0.190    0.082 0.112    0.078 0.101    0.554 0.764    0.017
1500 NBE     1.021 0.180    0.087 0.114    0.077 0.098    0.458 0.604    0.029
GMM     2.848 0.504    0.178 0.281    0.204 0.274    1.232 1.794    0.336
NRE     1.192 0.211    0.084 0.113    0.090 0.117    0.636 0.911    0.021
TRE     0.945 0.167    0.074 0.100    0.070 0.091    0.502 0.698    0.014
2000 NBE     0.909 0.161    0.079 0.103    0.072 0.091    0.416 0.552    0.025
Table 1: Estimation error comparison across sequence lengths $1000,1500,2000$ (top to bottom blocks) for GMM, NRE, TRE, and NBE (top to bottom rows within each block). From left to right, we display mean $L^{1}$ and $L^{2}$ distances between the true and inferred ACF functions, MAE and RMSE for the marginal parameters $\mu,\sigma$ and $\beta$, and mean KL divergence between true and inferred marginal distributions.
To confirm the superiority of TRE, we conduct a simulation study on $10^{4}$ samples $(\bm{x},\bm{\theta})$ drawn from the model. For each realization, we determine the point estimate via the BFGS optimization routine initialized at the true parameter $\bm{\theta}$, and summarize the estimation error in Table 1. We treat marginal and ACF parameters separately. For $(\mu,\sigma,\beta)$, we report the mean-absolute error (MAE), root mean-squared error (RMSE), and KL divergence between the true and inferred marginal distributions. For $\left(\gamma_{\textrm{acf}},\eta_{\textrm{acf}}\right)$, we directly compare the ACFs using the $L^{1}$ and $L^{2}$ distances rather than comparing individual parameters, as vastly different parameter values can sometimes yield nearly identical ACFs. While both TRE and NRE are trained on time series of length 1500, the LSTM encoder enables inference with time series of arbitrary length. Accross all metrics and lengths $k$, TRE substantially outperforms NRE and GMM. A natural question is how TRE compares to neural Bayes estimators (NBE), customized to target point summaries of the posterior distribution (Sainsbury-Dale et al. 2024). Training NBEs for trawl processes is tricky, as optimizing for the parameter-wise MAE or MSE fails due to the ACF weak-identifiability. We therefore train NBEs with multiple losses; see Section LABEL:sm:point_estimators for details. Table 1 shows the results for the NBE trained with $L^{2}$ ACF distance and MSE loss for marginal parameters. Even though NBEs have the advantage of directly optimizing the evaluation metric during training, we see that TRE achieves comparable or superior performance while providing the full posterior rather than just point estimates.
4.3 Calibration, coverage, and posterior checks
Although TRE point estimators perform well when the TRE model trained on time series of length $1500$ is applied to time series of other lengths, the encoder’s hidden state may encode length-specific features through its weights and biases. As we show below, the level sets of the likelihood approximation become distorted as the input length varies. To address this mismatch and amortize across different input lengths, we apply beta-calibration to multiple datasets $(\bm{x},\bm{\theta})$. Specifically, we take the models trained with $k=1500$ and calibrate them using datasets where the trawl process realizations $\bm{x}$ have lengths $k=1000$, $1500$ and $2000$. Importantly, calibration modifies the geometry of the likelihood approximation, but we confirm empirically that it does not change the estimation-errors from Table 1. We evaluate the accuracy of the resulting likelihood approximation both before and after calibration using the coverage diagnostics from Section 3.5.1 and metrics from Section 3.5.3.
Figure 4: Comparison of the coverage deviation $\mathcal{C}_{\alpha}-\alpha$ before and after beta-calibration. Positive values indicate underconfidence, while negative ones indicate overconfidence. Top row: NRE and TRE. Bottom row: component NREs within TRE. Beta-calibration yields near-perfect coverage for the TRE, consistently improving TRE performance across all lengths $1000$, $1500$ and $2000$.
Figure 4 displays the deviations $\mathcal{C}_{\alpha}-\alpha$, representing departures from perfect calibration across the trained classifiers. Starting with the bottom row of the figure, we show results for the ACF, $\beta$, $\mu$, and $\sigma$ classifiers used within the TRE. These classifiers vary in their degree of miscalibration and thus require different adjustments. As discussed in Section 3.1, the higher the value of $\mathcal{S}$, the harder it is to learn meaningful curves. Consistent with this, the ACF classifier requires the least calibration and also has the lowest $\mathcal{S}$ value (see Figure 2). Turning to the top row of the figure, we compare NRE and TRE. In the case of TRE, calibration is applied individually to each component classifier rather than to the combined estimator as a whole. This is because individual calibration addresses the distinct miscalibration levels across components, whereas calibrating the combined estimator applies a correction that fails to account for component-specific differences. Further, the combined classifier’s outputs are very close to 0 or $1$, and calibration becomes unreliable: see NRE for $k=1000$, where coverage degrades. Overall, the calibrated TRE has near perfect coverage, and outperforms both uncalibrated TRE and (uncalibrated or calibrated) NRE.
    NRE TRE ACF $\beta$ $\mu$ $\sigma$
    uncal cal    uncal cal    uncal cal    uncal cal    uncal cal    uncal cal
BCE     0.039 0.036    0.036 0.025    0.441 0.430    0.230 0.223    0.308 0.299    0.325 0.317
$\mathcal{S}$     5.455 5.232    6.038 6.177    0.967 0.995    2.026 2.064    1.576 1.615    1.468 1.504
1000 $\mathcal{B}$     0.987 1.000    0.978 0.999    0.960 1.000    0.968 0.999    0.961 1.000    0.964 0.999
$W$     0.018 0.124    0.119 0.015    0.044 0.003    0.043 0.006    0.052 0.003    0.040 0.006
ECE     —    —    0.034 0.008    0.019 0.003    0.026 0.004    0.024 0.005
BCE     0.023 0.022    0.015 0.015    0.388 0.388    0.199 0.199    0.268 0.268    0.285 0.285
$\mathcal{S}$     6.034 5.920    6.889 6.949    1.178 1.186    2.237 2.267    1.789 1.805    1.685 1.690
1500 $\mathcal{B}$     1.005 1.000    1.000 1.000    1.000 1.000    1.000 1.001    0.998 1.001    1.005 1.002
$W$     0.129 0.100    0.013 0.006    0.004 0.003    0.017 0.013    0.005 0.006    0.021 0.015
ECE     —    —    0.001 0.002    0.003 0.001    0.001 0.001    0.004 0.002
BCE     0.020 0.016    0.012 0.010    0.368 0.364    0.189 0.186    0.253 0.250    0.268 0.264
$\mathcal{S}$     6.289 6.374    7.298 7.442    1.276 1.309    2.338 2.389    1.895 1.926    1.788 1.818
2000 $\mathcal{B}$     1.009 1.000    1.004 1.000    1.021 1.001    1.015 1.001    1.016 1.001    1.022 1.001
$W$     0.205 0.097    0.106 0.030    0.042 0.013    0.064 0.033    0.036 0.010    0.059 0.022
ECE     —    —    0.021 0.004    0.012 0.002    0.012 0.001    0.017 0.001
Table 2: Comparison of NRE, TRE and (ACF, $\beta$, $\mu$, $\sigma$) TRE components (left to right) across different sequence lengths $1000$, $1500$, $2000$ (top to bottom blocks) based on the BCE, $\mathcal{S}$, $\mathcal{B}$, $W$ and ECE metrics (top to bottom rows within each block). Values are shown before (uncal) and after (cal) beta-calibration.
Next, Table 2 displays the BCE, $\mathcal{S}$, $\mathcal{B}$ and expected calibration error (ECE) metrics. We also report $W=\int_{0}^{1}\left|\mathcal{C}_{\alpha}-\alpha\right|\mathrm{d}\alpha$, which measures the overall deviation between the empirical and theoretical coverage. This metric corresponds to the Wasserstein distance between the distribution of posterior ranks and the uniform distribution, and can be interpreted as a rank check; see Section LABEL:SM:ranks. The calibrated TRE performs best across all metrics, with beta-calibration consistently improving individual classifier performance too. An interesting case is the NRE at $k=1000$, where the inherent underconfidence observed at $k=1500$ counterbalances, on average, the overconfidence gained from inputting shorter time series. However, the BCE and $\mathcal{S}$ metrics reveal that despite the reasonable coverage, the classifier’s quality is substantially worse than the corresponding TRE for $k=1000$.
Remark 4.1.
In this section, we employed beta-calibration to make a fair comparison between TRE and NRE. While isotonic regression is more flexible, it prevents gradient-based MCMC, thereby efficient posterior sampling and inference for the NRE. By contrast, the advocated Chebyshev polynomial approach is compatible with isotonic regression, and Section LABEL:SM:extended_beta_vs_iso shows that it can outperform beta-calibration for TRE. Methodologically, we note that the approximate densities $\theta^{i}\to\hat{p}(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})$ may not integrate exactly to $1$. Sequential sampling in the dimensions of $\bm{\theta}$ may thus yield different results to those obtained by combining the approximate ratios within an MCMC, as the NRE components’ weights differ. We do not observe significant deviations at this stage, and refer this for future research.
5 Application
We apply our methodology to daily energy demand data from January 1, 2019, to December 31, 2024, across nine U.S. regions, reported by respondents with acronyms AZPS, BPAT, CISO, DUK, ERCO, FPL, MISO, NYIS, and PJM. The computer code used to download the datasets and perform the analysis is provided alongside the paper. To account for non-stationary behaviour, the data are deseasonalised using the LOESS algorithm, as implemented in Python’s statsmodels package, version 0.14.5. Figure 5 shows the decomposition of the time series from Arizona Public Service Company (AZPS) into two parts: a non-stationary component, which captures the trend and seasonal fluctuations, and a stationary component, represented by the residuals after removing trend and seasonality.
We apply the TRE methodology using the pre-trained posterior to infer the parameters of the trawl process to the residuals for all nine time series and display results for AZPS in Figure 6. We also summarize results for all time series in Table 3. The amortized TRE posterior sampling based on Chebyshev polynomial approximations is highly efficient: generating $10^{3}$ independent samples takes under one second on a single Intel i9-14900K (3.20 GHz) CPU core, with further acceleration possible on a GPU, enabling real-time inference. Another advantage is that we avoid costly gradient-descent runs with multiple initializations to accurately determine the MAP, since the maximum over the posterior samples already provides a reliable starting point.
Although the trawl process models the data reasonably well, there is some slight lack of fit at short time lags resulting from the inability to fully remove seasonality, which can still be observed in the right part of Figure 6. A more elegant approach is to incorporate non-stationarity in the trawl model. Periodic trawl processes (Veraart 2024), which introduce seasonal effects through a kernel, provide a natural extension. Moreover, by indexing the sets $A_{t}(\mathbf{z})$ in both time $t$ and space $\mathbf{z}$, this framework generalizes seamlessly to parsimonious spatio-temporal models where $(t,\mathbf{z})\to L(A_{t}(\mathbf{z}))$, allowing joint analysis of temperature data from multiple stations rather than treating them as independent series, while preserving fast simulation algorithms and the TRE posterior inference framework.
Figure 5: Arizona electricity demand in megawatt-hours as reported by Arizona Public Service Company (AZPS). Left: original demand time series together with estimated trend and seasonality; right: residuals.
Figure 6: Histogram, kernel density estimate (KDE), and the NIG marginal distribution corresponding to the MAP-fitted parameters of the trawl process. Right: Empirical autocorrelation function (ACF) with $95\%$ bootstrapped confidence intervals (CI), the MAP-inferred ACF corresponding to the MAP-fitted parameters with its $95\%$ theoretical CI, and the posterior mean and median ACFs. The posterior mean and median ACFs are computed by taking the lag-wise mean and median of the ACFs evaluated at parameter draws from the posterior distribution of the trawl process parameters. Bartlett’s CI are also shown for reference.
Dataset $\mu$ $\sigma$ $\beta$ ACF(1) $\tau_{\text{eff}}$
AZPS -36 (-1130, 1055) 6184 (5696, 6866) 0.041 (-0.131, 0.214) 0.880 (0.856, 0.902) 25
ERCO -572 (-13412, 12261) 78957 (75485, 88614) 0.011 (-0.214, 0.194) 0.849 (0.826, 0.873) 19
CISO 186 (-4684, 5064) 28915 (27311, 32522) 0.087 (-0.084, 0.234) 0.856 (0.832, 0.878) 20
NYIS 704 (-3059, 3876) 23899 (22538, 26332) 0.058 (-0.133, 0.263) 0.832 (0.822, 0.846) 17
PJM -2965 (-21502, 18128) 130295 (124304, 145757) 0.071 (-0.093, 0.288) 0.842 (0.825, 0.861) 17
MISO -2190 (-16929, 11288) 86804 (81161, 95947) 0.072 (-0.115, 0.222) 0.862 (0.839, 0.883) 21
FPL 120 (-4433, 2980) 24955 (24000, 28389) -0.111 (-0.378, 0.085) 0.838 (0.823, 0.855) 17
DUK -209 (-3137, 3472) 23136 (21866, 25014) 0.079 (-0.149, 0.254) 0.830 (0.821, 0.845) 17
BPAT 215 (-1201, 1581) 8082 (7472, 9178) 0.144 (-0.038, 0.332) 0.868 (0.845, 0.891) 22
Table 3: MAP estimates of the trawl process parameters (columns): $\mu,\sigma$, $\beta$, as well as the lag-11 correlation and the effective correlation range, $\tau_{\text{eff}}$, defined as the lag at which the MAP-inferred ACF falls below $0.05$, across nine energy-demand time series (rows). Numbers in brackets are $95\%$ confidence intervals.
6 Conclusion
In this paper, we advance the existing neural ratio estimation methodology for amortized simulation-based inference (SBI) along four directions. First, we propose telescoping ratio estimation (TRE) in the SBI setup and decompose the global classification task into multiple subtasks, thereby improving training efficiency and mitigating the “curse of parameter space dimensionality”. Learning parameters sequentially makes the estimation order a key design choice, and we provide guidance on how to select it. Second, we propose a novel, MCMC-free sequential sampling approach that is GPU-compatible and that leverages Chebyshev polynomial approximations for efficient inverse sampling. Third, we develop novel, granular and nearly instantaneous diagnostic checks on the quality of the learnt likelihood ratio. Beyond faster inference, it also makes large-scale model comparison practically feasible. Fourth, we show that post-training calibration can be integrated into TRE to both improve the learnt ratio and further extend amortization, for instance accross different lengths of the input time series. Through extensive simulations and an application to energy demand data based on non-Gaussian trawl process models, we have demonstrated significant improvements over existing methods and performance on par with neural Bayes estimators.
Our work also opens further research avenues. The methodology can be extended from temporal to spatio-temporal settings such as ambit fields with only network architecture modifications—a development that would be novel for such fields. Additionally, amortization could encompass not only varying sequence lengths but also irregular spacing and partial observations while preserving the reliability of the likelihood estimates. Finally, future work could also benchmark our approach against other SBI methods to better understand its comparative advantages.
Acknowledgments The authors would like to thank Hugo Chu and Maja Schneider for helpful discussions and comments on the manuscript. The dataset used in Section 5 is available at https://www.eia.gov/opendata/. Computer code is available at https://github.com/danleonte/Simulation-based-inference-via-telescoping-ratio-estimation-for-trawl-processes.